OpenStudy (anonymous):

GEOMETRY!!! Whats an equation of the line parallel to the given line that contains C? C(6,0); y=1/3x

OpenStudy (kmeis002):

Given two lines: \[y_1 = m_1x+b_1 \\ y_2 = m_2x+b_2\] Do you know what makes those two lines parallel? (any two lines that is)

OpenStudy (anonymous):

plug into y=mx+b and solve, then take what you got for b and take the current slope you have and make an equation

OpenStudy (anonymous):

Nope, i have no clue. My teachers an idiot and wont teach crap.

OpenStudy (kmeis002):

Slope is usually denoted by the constant (m) in linear equations and is given by: \[m \frac{y_2-y_1}{x_2-x_1} \] Slope shows the change in y with respect to the change in x (if x changes 1 unit, how much will y change). It refers to how steep a line is: Two lines are parallel if their slopes are the same. So for example: \[ y_1 = 2x+5\\ y_2 = 2x - 2 \] These two lines are parallel because they both have a slope of 2, but different y intercepts. The general form of a line is given by: \[ y = \underbrace{m}_{\text{slope}} x + \underbrace{b}_{\text{y-intercept}}\]. The b tells you where the line interesects the y axis. Now for your problem. We know the slope of the first line is 1/3. So we know our new slope must be the same. We get the partial equation: \[y = \frac{1}{3}x +b \] We also know the line passes through (6,0). THis will help us solve for b. Plug in x=6 and y=0 into our equation and we get: \[0 = \frac{1}{3}6 + b = 2+b \quad \to \quad b=-2\] Solving for b gives us -2, we can then finally write our new line: \[y = \frac{1}{3}x -2 \]